OpenStudy (anonymous):

(2/x+1)-(1/x-2)=-1/2 Show steps if possible thank you(:

OpenStudy (anonymous):

lol easy :D \[2/(x+1) - (1/x-2) = -1/2\] \[[[2(x-2)] - [1(x+1)]]/[(x+1)(x-2)]\] \[(2x-4 - x-1) / (x+1)(x-2)\] \[x-5 / (x+1)(x-2) = -1/2\] now just convert into quadratic and then solve for x

OpenStudy (anonymous):

How do i do that? i'm not very smart when it comes to this type of math!

OpenStudy (anonymous):

not doing your work for you. either ask for help (teacher) or learn to use google http://www.purplemath.com/modules/solvquad.htm

OpenStudy (anonymous):

multiply the x+1 and x-2 to the -1/2

OpenStudy (anonymous):

tutordiffeq, you have to let them do at least a part or they gain nothing

OpenStudy (anonymous):

i didnt give the answer i just showed the way

OpenStudy (anonymous):

that is a confusing place to leave them. once she gets away from the fractions, she should be good.

OpenStudy (anonymous):

thank you; tutor that's where i get stuck is the fractions; the thing i'm confused on is; the x-5 on top; so i was wondering what to do.

OpenStudy (anonymous):

Okay i got x-5/x^2-x-2=-1/2; i'm confused what to do next.!

OpenStudy (anonymous):

move the bottom part of the fraction to the other side of the = by multiplication

OpenStudy (anonymous):

so multiply -1/2(x^2-x-2)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks. :) should the answers be -4,3?

OpenStudy (anonymous):

-1/2x^2-1/2x+6=0 use quadratic and you figure it out. no calc. here for me to use.

OpenStudy (anonymous):

Thank you so much for your help; i've learned the quadratic formulas and everything; just not problems like this.!

OpenStudy (anonymous):

sometimes you have to think what form can you deal with it in and then look at how you can get it there.